# Learned Contextual LiDAR Informed Visual Search in Unseen Environments

### 28 Feb 2024 v3

@@ I INTRODUCTION
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. Use of scans is often limited to obstacle avoidance, localization and mapping. In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information. At the same time, the potential to use additional information from LiDAR scans for online control has grown with access to parallel computing and more sophisticated GPUs. While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], ref. [17] notes these methods are limited by the high computational requirements to handle to large input size of 3D cloud data. In contrast, real time projection-based methods aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26]. There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31]. Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [3, 14, 36, 42], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [43]. There exists a gap of visual search planners utilizing additional information available in LiDAR scans from segmentation. This work combines these concepts to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
@@ III-E2 Viewpoint Selection
  • Subtract path distance from robot position to the viewpoint
  • Add a reward for expected number of unknown cells discovered at that viewpoint
  • Add a reward for viewpoints with frontiers near the path to the candidate viewpoint
  • Add a reward for viewpoints that result in inspection of a scaled number non-map points
@@ IV-B Simulation and Baseline Comparison
Figure 4: The Apartment simulation environment (top row) in the Easy configuration (a) and Hard Configuration (b). The Office (c) shows the robot starting position as a red arrow. Easy setup involves the search target to the left of the cyan line and Hard involves the search target on the right of the cyan line.
Two Gazebo world environments are tested with easy and hard setups in each. The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES (see Fig. 4(a-b)). Here the easy setting incorporates fewer objects in the environment while hard is defined by more objects in the environment. The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position (see Fig. 4(c)). The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
@@ IV-C Hardware Experiments
Experiments take place in a 20mx30m indoor apartment setting and a 15mx38m hallway setting, shown in Fig. 5. The environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy. A Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera is utilized. The search target of interest is one or two small rolling suitcases and YOLOv5 [18] is used for detection of the object. Objects are localized by first converting PointCloud data into the optical frame and then by conversion to pixel coordinates using camera intrinsics. A time lapse of one experiment in the apartment is shown in Fig. 6 where the search target is situated behind the robot starting position. This example demonstrates the benefit of non-map information. In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the one on the left hand side. However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby. Finally, we compare the performance of the trained classifier in these out of distribution environments to ground truth information. The scan classifier accuracy is also evaluated in the two out of distribution experimental environments. The results show $84.0168\%$ accuracy in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment. These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments. The success of LIVES onboard a real robot verify hypothesis (3). Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home.
@@ V Conclusions
We present a method for visual search in unseen environments using contextual LiDAR information to inform planning. Results show LIVES to outperform the closest of several baseline methods in varied environments by 10-30%. The proposed approach is deployed on a real robot in an unseen environment to certify the learning method and visual planner for real-world scenarios. Ongoing work aims to extend the LIVES planner to multiple robots.

### 29 Feb 2024 v4

@@ I INTRODUCTION
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. Use of scans is often limited to obstacle avoidance, localization and mapping. In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information. At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing. While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], these methods are limited by the high computational requirements to handle 3D cloud data [17]. In contrast, real time projection-based methods aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26]. There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31]. Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [3, 14, 36, 42], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [43]. There exists a gap of visual search planners utilizing additional information available in LiDAR scans from segmentation. This work combines these concepts to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
@@ III-E2 Viewpoint Selection
  • Subtract path distance from robot position to viewpoint
  • Reward expected number of unknown cells discovered
  • Reward viewpoints with frontiers near the path to the candidate viewpoint
  • Reward viewpoints that result in inspection of a scaled number non-map points
@@ IV-B Simulation and Baseline Comparison
Two Gazebo world environments are tested with easy and hard setups in each. A figure depicting them is shown on the project website. The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES. The easy setting incorporates fewer objects in the environment while hard has more objects in the environment. The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position. The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
@@ IV-C Hardware Experiments
Experiments take place in a 20mx30m indoor apartment setting and a 15mx38m hallway setting, shown in Fig. 4. The environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy. A Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera is utilized. The search target of interest is one or two small rolling suitcases and YOLOv5 [18] is used for detection of the object. Objects are localized by first converting PointCloud data into the optical frame and then by conversion to pixel coordinates using camera intrinsics. A time lapse of one experiment in the apartment is shown in Fig. 5 where the search target is situated behind the robot starting position. This example demonstrates the benefit of non-map information. In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the one on the left hand side. However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby. Finally, we compare the accuracy of the trained classifier in these out of distribution environments to ground truth information. The evaluated accuracy in the two out of distribution experimental environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment. These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments. The success of LIVES onboard a real robot verify hypothesis (3). Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home.
@@ V Conclusions
We present a method for search in unseen environments using contextual information to inform planning. Results show LIVES to outperform several baseline methods in varied environments by 10-30%. Ablative studies on the map-free scan classifier are performed to validate the choice in architecture and training scheme. The proposed approach is deployed in an unseen environment to certify the learning method and visual planner for real-world scenarios. Ongoing work aims to extend the LIVES planner to multiple robots.